A pasture is twice as long as it is wide. Its area is 115,200 ft2. How wide is the pasture?
Answer by rfer(16322)

2x*x=115200
2x^2=115200
x^2=115200/2
x^2=57600
x=sqrt 57600
x=240 wide
2x=480 long